What value of x makes this equation true?
2 (x-4)-12=6 (x-2)+32

Answer by fractalier(6550)

Okay, from
2(x-4) - 12 = 6(x-2) + 32
2x - 8 - 12 = 6x - 12 + 32
2x - 20 = 6x + 20
Now add 20 and subtract 6x from both sides and get
-4x = 40
Now divide by -4 and get
x = -10